OpenStudy (litzyceron):

What is the surface area of this rectangular prism? A. 198 m2 B. 81 m2 C. 99 m2 D. 162 m2

OpenStudy (litzyceron):

i need help somebody plz :/

OpenStudy (litzyceron):

@MARC_

OpenStudy (litzyceron):

@ganeshie8

OpenStudy (litzyceron):

@Nerdsarecool

jimthompson5910 (jim_thompson5910):

see attached

OpenStudy (anonymous):

Surface Area of a Rectangular Prism = 2ab + 2bc + 2ac (a, b, and c are the lengths of the 3 sides) In words, the surface area of a rectangular prism is the area of the six rectangles that cover it. But we don't have to figure out all six because we know that the top and bottom are the same, the front and back are the same, and the left and right sides are the same. The area of the top and bottom (side lengths a and c) = a*c. Since there are two of them, you get 2ac. The front and back have side lengths of b and c. The area of one of them is b*c, and there are two of them, so the surface area of those two is 2bc. The left and right side have side lengths of a and b, so the surface area of one of them is a*b. Again, there are two of them, so their combined surface area is 2ab.

OpenStudy (litzyceron):

so is it 162 not sure ?

OpenStudy (litzyceron):

im not sure tho??

jimthompson5910 (jim_thompson5910):

SA = 2*(LW + WH + LH) L = 9 W = 6 H = 3 SA = 2*(LW + WH + LH) SA = 2*(9*6 + 6*3 + 9*3) SA = ???

OpenStudy (anonymous):

based on @jim_thompson5910 working\[SA=2(54+18+27)\]\[SA=?\]solve it @litzyceron

OpenStudy (litzyceron):

99

OpenStudy (anonymous):

don't forget to multiply it by 2

OpenStudy (anonymous):

\[SA=2(99)\]\[SA=?\]

OpenStudy (litzyceron):

so its a 198

jimthompson5910 (jim_thompson5910):

yep, 198

OpenStudy (anonymous):

yes,it is :)

OpenStudy (litzyceron):

ok thank you all for helping me @MARC_ & @jim_thompson5910

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (litzyceron):

i have one more question ?

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (litzyceron):

What is the surface area of this prism? A. 3312 m2 B. 3240 m2 C. 6480 m2 D. 1656 m2

jimthompson5910 (jim_thompson5910):

you'll need to post the pic of it

OpenStudy (litzyceron):

hold on

jimthompson5910 (jim_thompson5910):

what is the area of each triangle? hint: A = b*h/2

OpenStudy (litzyceron):

108 for the first one and the second one is 37.5 ?

jimthompson5910 (jim_thompson5910):

both triangles are congruent, so each one is 108

OpenStudy (litzyceron):

ohhhh ok

jimthompson5910 (jim_thompson5910):

the side panels are rectangles with length 30 and width 15 there are 3 side panels

jimthompson5910 (jim_thompson5910):

actually hold on

OpenStudy (litzyceron):

ok

jimthompson5910 (jim_thompson5910):

the two top side panels are rectangles with length 30 and width 15 the bottom rectangle is a rectangle with length 30 and width 18

OpenStudy (litzyceron):

ok

OpenStudy (litzyceron):

???

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):